# biblio.bib

### v1

@@ Acknowledgements
The clustering coefficient is instead a measure of the nodes’ tendency to congregate. While there are many possible definitions, we refer to the one given by Watts and Strogatz in [watts1998smallworld] since it is defined for the class of graphs we consider in this work, i.e. small-world networks. Specifically, it refers to the global clustering coefficient computed as average of the local ones. Given a node $q$ with $k_{q}$ neighbors (in other words it represents the degree of the site), we define the local clustering coefficient $C_{q}$ as the proportion $t_{q}$ of the number of triangles, i.e. the closed loops consisting of 3 nodes, containing $q$, counted with respect to the maximum number of allowed connections $k_{q}(k_{q}-1)/2$ (if $k_{q}$ is greater than 1): the latter is also referred in the literature as the overall amount of triplets, or subgraphs composed of three nodes connected with at least two edges. Using the properties of the graph’s adjacency matrix $A$ ($A_{ij}=1$ if vertices $i$ and $j$ are connected, 0 otherwise), specifically the one that says that its powers indicate the number of walks between two sites, one can compute $t_{q}$ as the matrix element $(A^{3})_{qq}$ divided by 2 (graphs are undirected, so this allows to count the unique triangles). In summary, the clustering coefficient $\mathcal{C}$ comes as follows:

### v2

@@ Appendix C Characterizing complex graphs: average distance and clustering
The clustering coefficient is instead a measure of the nodes’ tendency to congregate. While there are many possible definitions, we refer to the one given by Watts and Strogatz in [78] since it is defined for the class of graphs we consider in this work, i.e. small-world networks. Specifically, it refers to the global clustering coefficient computed as average of the local ones. Given a node $q$ with $k_{q}$ neighbors (in other words it represents the degree of the site), we define the local clustering coefficient $C_{q}$ as the proportion $t_{q}$ of the number of triangles, i.e. the closed loops consisting of 3 nodes, containing $q$, counted with respect to the maximum number of allowed connections $k_{q}(k_{q}-1)/2$ (if $k_{q}$ is greater than 1): the latter is also referred in the literature as the overall amount of triplets, or subgraphs composed of three nodes connected with at least two edges. Using the properties of the graph’s adjacency matrix $A$ ($A_{ij}=1$ if vertices $i$ and $j$ are connected, 0 otherwise), specifically the one that says that its powers indicate the number of walks between two sites, one can compute $t_{q}$ as the matrix element $(A^{3})_{qq}$ divided by 2 (graphs are undirected, so this allows to count the unique triangles). In summary, the clustering coefficient $\mathcal{C}$ comes as follows: